A particular engine has a power output of 6.00 kW and an efficiency of 22.0%. Assume the engine expels 7000 J of energy in each cycle. (a) Find the energy absorbed in each cycle. kJ (b) Find the time required to complete each cycle. s A particular engine has a power output of 6.00 kW and an efficiency of 22.0%. Assume the engine expels 7000 J of energy in each cycle.

(a) Find the energy absorbed in each cycle. kJ

(b) Find the time required to complete each cycle. s
Answer #1

a) 0.22 = 1 - (Qc/Qh)

Qc/Qh = 1 - 0.22

Qc/Qh = 0.78

Qh = 7000 / 0.78

= 8974 J = 8.974 KJ

b) W = Qh - Qc = 8974 - 7000 = 1974 J

P = W / delta t

delta t = 1974 / 6000

= 0.329 s